calculate the vapour pressure lowering form by addition of 100g of sucrose to 1000g of water if the vapour pressure of pure water is 23.8 mmhg

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
P-Ps/P = X solute = n/(n +N)
n = 100/342 = 0.292
N = 1000/18 = 55.5
P = 23.8 mm Hg
Hence
Delta P /23.8 = 0.292/(0.292 + 55.5)
Delta P = 0.125 mm Hg
Regards
Arun (askIITians forum expert)